# Spin-operator form factors of the critical Ising chain and their finite volume scaling limits

### v1

@@ Abstract
In this work, we provide a self-contained derivation of the spin-operator matrix elements in the fermionic basis, for the critical Ising chain at a generic system length $N\in 2Z_{\geq 2}$. The approach relies on the near-Cauchy property of certain matrices formed by the Toeplitz symbol in the critical model, and leads to simpler product formulas for the dressing functions in terms of square root functions. These square root products allow fully dis-singularized integral representations. In the finite volume scaling limit, they further reduce to the Binet’s second integral for the gamma function logarithm and its Hermite’s generalization. As such, all the matrix elements in the scaling limit allow simple product formulas in terms of the gamma function at integer and half-integer arguments, and are rational numbers up to $\sqrt{2}$. They are exactly the spin-operator form factors of the Ising CFT in the fermionic basis, whose explicit forms are much less well known in comparison to the finite-volume form factors in the massive theory. We also fully determine the normalization factor of the spin-operator and show explicitly how the coefficient $G(\frac{1}{2})G(\frac{3}{2})$ appear through a ground state overlap. As a by-product, we obtain a Mellin-Barnes representation of the Glaisher constant with two $\zeta$ functions
@@
whose arguments are symmetric under $u\rightarrow-u$.
@@ 1 The critical Ising chain, square-root products and main results
The finite volume spin-operator form factors of the massive fermion Ising-QFT Fonseca:2001dc were well known, but what about their massless limits? Will the results become meaningless, or will they reduce to the spin-operator matrix elements of the Ising CFT in the fermionic basis, that were computed only recursively in Yurov:1991my? To the author’s knowledge, it seems that many references who used the massive form factors in Fonseca:2001dc, including this paper itself, forgot to mention their convergence to a nontrivial finite volume massless scaling limit. Instead, the integrals in the leg functions were usually written in certain forms, creating impressions of their ill-posedness or triviality at the zero fermion mass [^1] For example, in Ref. Gabai:2019ryw, the leg function was written with an overall $m^{2}$. Still, it is easier to spot a non-trivial massless limit from the representation in Gabai:2019ryw than the rapidity space integral in Fonseca:2001dc..
@@
Similarly, in the literature, there are plenty discussions and derivations Lisovy:2001gc; Bugrij:2003xhk; Bugrij:2004sac; Gehlen_2008; Iorgov:2010yv; Iorgov:2010ie for the finite volume form factors in the Ising model and the Ising chain. These works are very non-trivial, but most of them were assuming $T>T_{c}$ or $T<T_{c}$, and forgot to mention what happened at the critical point. For example, in the published version of Gehlen_2008, the authors said explicitly after the Eq. (78) that the results are valid both when $g>1$ and $0\leq g<1$, but just omitted the critical point $g=1$. Very often, the formulas were written with a refactor such $\xi=(g^{2}-1)^{\frac{1}{4}}$ in front, a factor that vanishes at the critical point $g=1$, like the expressions given to the leg functions in the massive theory, despite that in the proper way of taking the $g\rightarrow 1^{+}$ limit, the results are actually finite and as we will show, lead exactly to the form factors at the critical point.
@@
Meanwhile, in the literature, there were also not too many discussions on the spin-operator form factors in the Ising-CFT, despite that they are clearly universal, beautiful and unique quantities: the massless fermion basis of the Ising CFT is perhaps among the few integrable massless scattering basis Bazhanov:1994ft; Bazhanov:1996aq of 2D CFTs, that allow explicit expressions for local operator’s matrix elements between them. On the other hand, in comparison to the Virasoro basis, it could be the massless scattering basis, that respect better the scale-full deformations of CFTs either to the integrable lattice models in the UV Boos:2010qii, or to the massive QFTs in the IR. It is possible, that for certain non-inclusive high-energy limits that are more sensitive to the structures of the Hilbert space, one needs not just the standard CFT perturbation theory for point-like objects, but also the correct massless basis, to write out the correct high-energy asymptotics. These questions deserve more investigations, and were among the motivations of this work. To the author’s opinion, CFTs would not be universal, if they could not be reached.
@@
Given all above, the author believes that it is reasonable and necessary to provide a self-contained derivation for the spin-operator form factors in the 2D Ising chain, directly at the critical point. Below we will provide such a derivation, and show how in the finite volume scaling limit, the results converge to the Ising CFT versions. We believe that the products, integral representations and the large $N$ asymptotics in this work are simple yet beautiful. We also believe that the explicit formulas provided by this work could be of practical usages (such as the TFCSA applications Yurov:1991my).
@@ 1.2 The critical Ising chain
After introducing the sets $S_{\pm}$, we set up the conventions for the critical Ising chain. We choose the ,+-,+ convention of the Hamiltonian
@@
$\lambda_{2k}(\beta)=\beta^{k}u(\beta)\ ,\ \lambda_{2k+1}(\beta)=\beta^{k}\nu(\beta)\ ,\ \beta^{N}=-1\ ,$ (18)
@@ 1.3 Square-root products and form-factors at finite NN
The advantages of using the rescaled spin operator are two folds. First, computation of the rescaled form factors is naturally separated from the computation of the ground-state overlap in our algorithm. Second, it is the rescaled form-factors that allow finite scaling limit, defined as
@@
$N\rightarrow\infty,z\rightarrow 1,\ N(z-1)\rightarrow\pi i\times(2m+1)\ ,\ m\in Z\ .$ (56)
Clearly, the numbers $n$ and $m+\frac{1}{2}$ correspond to the quantization levels on a cylinder. The vacuum overlap, on the other hand, approaches to zero at a power-law exactly specified by the scaling dimension of the spin operator. We will show that
@@
$S^{2N}_{+}=\{\sqrt{\beta};\beta\in S_{+}\}\cup\{-\sqrt{\beta};\beta\in S_{-}\}\ ,$ (66)
@@ 1.4 Scaling limits of the form factors
$\rightarrow\int_{0}^{\infty}\frac{dt}{\pi}\frac{2\pi|\alpha|}{t^{2}+\alpha^{2}}\ln\frac{1-e^{-t}}{1+e^{-t}}-\frac{\pi|\alpha|}{96N^{2}}+{\cal O}\left(\frac{\pi|\alpha|^{3}}{N^{4}}\right)\ .$ (95)
@@
$\mathfrak{b}(\alpha)=\int_{0}^{\infty}\frac{dt}{\pi}\frac{\alpha}{t^{2}+\alpha^{2}}\ln\frac{1-e^{-t}}{1+e^{-t}}\equiv\ln\sqrt{\alpha}+\ln\frac{\Gamma\left(\alpha+\frac{1}{2}\right)}{\sqrt{2\pi}\Gamma(\alpha+1)}\ ,\ \alpha>0\ .$ (96)
@@ 2.2 The Wick contraction
${\cal M}_{2k+1,2l}(\beta_{i},z_{j})={\cal M}_{2k+1,2l}=(-i)^{l}\langle\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}\rangle\ .$ (165)
The extra factor of $(-i)^{l}$ is equivalent to the redefinition
@@
$(-i)^{l}\langle\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}\rangle={\cal M}_{2k+1,2l}(\beta_{i},z_{j})$
@@ 2.3 The vacuum overlap
where $A$ is the Glaisher constant. On the other hand, the term with $\tanh^{2}\frac{t}{4N}$ represents the sub-leading corrections and can be easily expanded to an arbitrary order in $\frac{1}{N}$
@@
$\sum_{i=1}^{n}\ln i!=(n+1)\ln n!-\ln\prod_{i=1}^{n}i^{i}=(n+1)\ln n-\ln H(n)\ ,$ (210)
$\ln H(n)\rightarrow\left(\frac{n^{2}+n}{2}+\frac{1}{12}\right)\ln n-\frac{n!}{4}+\ln A\ .$ (211)
@@ 2.4 Scaling limit of the two-point correlator
This is exactly the standard CFT two-point correlator. As such, after multiplying back the factor $|\langle\Omega_{R}|\Omega_{NS}\rangle|^{2}$, at small $r\ll L$, one has then
@@ 3 Comparison with literature and further comments
$\kappa(\mu L,\mu L)=\frac{1}{\pi}\int_{1}^{\infty}\frac{dt}{t\sqrt{t^{2}-1}}\ln\frac{1-e^{-\mu Lt}}{1+e^{-\mu Lt}}=\frac{1}{2}\ln mL+{o}(1)\ ,$ (235)
@@
again consistent with the $d(1)=1$. Two particle functions clearly simplify to the $\Pi$ factors in Eq. (104) and Eq. (105). As such, up to the phase factors and the overall normalization convention, the massless limit of the finite volume massive matrix elements again match the finite volume scaling limit of a critical lattice model.
@@
where we have used the relation p(1)q(1)=2ip(1)q(1)=-2i in Eq. (71). As such, all the form factor squared in Gehlen_2008; Iorgov:2010yv, in the proper $g\rightarrow 1^{+}$ limit approached from the disordered side, agree with our results derived directly at the critical point.
@@
where $K(w)$ is the complete Elliptic integral of the second Kind. The $\frac{K(w)}{\pi}$ is exactly half of the one particle form factor contribution $f^{(1)}_{0,0}(t)$ to the diagonal correlator at zero separation Boukraa:2006bt. If such relations could be established at high orders, then the identity Eq. (245) becomes understandable: $\sigma_{00}^{2}=1$, while the form factor expansion in Boukraa:2006bt contains an overall $(1-t)^{\frac{1}{4}}$. This implies that all the form factors at zero separation must sum to $(1-t)^{-\frac{1}{4}}$. It is also possible that differential equations in $w$ could be established, similar to the Painleve VI for the parameter derivatives 10.3792/pjaa.56.405.
@@ Acknowledgements.
The author thanks Zoltan. Bajnok for discussions and notification of the references Yurov:1991my; Fonseca:2001dc.

### v2

@@ Abstract
In this work, we provide a self-contained derivation of the spin-operator matrix elements in the fermionic basis, for the critical periodic Ising chain at a generic system length $N\in 2Z_{\geq 2}$. The approach relies on the near-Cauchy property of certain matrices formed by the Toeplitz symbol in the critical model, and leads to simpler product formulas for the dressing functions in terms of square root functions. These square root products allow fully dis-singularized integral representations. In the finite volume scaling limit, they further reduce to the Binet’s second integral for the gamma function logarithm and its Hermite’s generalization. As such, all the matrix elements in the scaling limit allow simple product formulas in terms of the gamma function at integer and half-integer arguments, and are rational numbers up to $\sqrt{2}$. They are exactly the spin-operator form factors of the Ising CFT in the fermionic basis, whose explicit forms are much less well known in comparison to the finite-volume form factors in the massive theory. We also fully determine the normalization factor of the spin-operator and show explicitly how the coefficient $G(\frac{1}{2})G(\frac{3}{2})$ appear through a ground state overlap. As a by-product, we obtain a Mellin-Barnes representation of the Glaisher constant with two $\zeta$ functions
@@
which allows to establish analytically the small mass asymptotics of a related crossover function.
@@ 1 The critical Ising chain, square-root products and main results
The finite volume spin-operator form factors of the massive fermion Ising-QFT Fonseca:2001dc were well known, but what about their massless limits? Will the results become meaningless, or will they reduce to the spin-operator matrix elements of the Ising CFT in the fermionic basis, that were computed only iteratively in Yurov:1991my? To the author’s knowledge, it seems that many references who used the massive form factors in Fonseca:2001dc, including this paper itself, forgot to mention their convergence to a nontrivial finite volume massless scaling limit. Instead, the integrals in the leg functions were usually written in certain forms, creating impressions of their ill-posedness or triviality at the zero fermion mass [^1] For example, in Ref. Gabai:2019ryw, the leg function was written with an overall $m^{2}$. Still, it is easier to spot a non-trivial massless limit from the representation in Gabai:2019ryw than the rapidity space integral in Fonseca:2001dc..
@@
Similarly, in the literature, there are plenty discussions and derivations Lisovy:2001gc; Bugrij:2003xhk; Bugrij:2004sac; Gehlen_2008; Iorgov:2010yv; Iorgov:2010ie for the finite volume form factors in the Ising model and the Ising chain. These works are very non-trivial, but most of them were assuming $T>T_{c}$ or $T<T_{c}$, and forgot to mention what happened at the critical point. For example, in the published version of Gehlen_2008, the authors said explicitly after the Eq. (78) that the results are valid both when $g>1$ and $0\leq g<1$, but just omitted the critical point $g=1$. Very often, the formulas were written with a refactor such as $\xi=(g^{2}-1)^{\frac{1}{4}}$ in front, a factor that vanishes at the critical point $g=1$, like the expressions given to the leg functions in the massive theory, despite that in the proper way of taking the $g\rightarrow 1^{+}$ limit, the results are actually finite and as we will show, lead exactly to the form factors at the critical point.
@@
Meanwhile, in the literature, there were also not too many discussions on the spin-operator form factors in the Ising-CFT, despite that they are clearly universal, beautiful and unique quantities: the massless fermion basis of the Ising CFT is perhaps among the few integrable massless scattering basis Bazhanov:1994ft; Bazhanov:1996aq of 2D CFTs, that allow explicit expressions for local operator’s matrix elements between them. On the other hand, in comparison to the Virasoro basis, it could be the massless scattering basis, that respect better the scale-full deformations of CFTs either to the integrable lattice models in the UV Boos:2010qii, or to the massive QFTs in the IR. It is possible, that for certain non-inclusive high-energy limits that are more sensitive to the structures of the Hilbert space, one needs not just the standard CFT perturbation theory for point-like objects, but also the correct massless basis, to write out the correct high-energy asymptotics. These questions deserve more investigations, and were among the motivations of this work.
@@
Given all above, the author believes that it is reasonable and necessary to provide a self-contained derivation for the spin-operator form factors in the 2D periodic Ising chain, directly at the critical point. Below we will provide such a derivation, and show how in the finite volume scaling limit, the results converge to the Ising CFT versions. We believe that the products, integral representations and the large $N$ asymptotics in this work are simple yet beautiful. We also believe that the explicit formulas provided by this work could be of practical usages (such as the TFCSA applications Yurov:1991my).
@@ 1.2 The critical Ising chain
After introducing the sets $S_{\pm}$, we set up the conventions for the critical periodic Ising chain. We choose the ,+-,+ convention of the Hamiltonian
@@
$\lambda_{2k}(\beta)=\beta^{k}u(\beta)\ ,\ \lambda_{2k+1}(\beta)=\beta^{k}\nu(\beta)\ ,\ \beta^{N}=+1\ ,$ (18)
@@ 1.3 Square-root products and form-factors at finite NN
The advantages of using the rescaled spin operator are two folds. First, computation of the rescaled form factors is naturally separated from the computation of the ground-state overlap in our algorithm. Second, it is the rescaled form-factors that allow finite-volume scaling limit, defined as
@@
$N\rightarrow\infty,z\rightarrow 1,\ N(z-1)\rightarrow 2\pi i\times m\ ,\ m\in Z+\frac{1}{2}\ .$ (56)
Clearly, the numbers $n\in Z$ and $m\in Z+\frac{1}{2}$ correspond to the quantization levels on a cylinder. The vacuum overlap, on the other hand, approaches to zero at a power-law exactly specified by the scaling dimension of the spin operator. We will show that
@@
$S^{2N}_{+}=\{\sqrt{\beta};\beta\in S_{+}\}\cup\{-\sqrt{\beta};\beta\in S_{+}\}\ ,$ (66)
@@ 1.4 Scaling limits of the form factors
$\rightarrow\int_{0}^{\infty}\frac{dt}{\pi}\frac{|\alpha|}{t^{2}+\alpha^{2}}\ln\frac{1-e^{-t}}{1+e^{-t}}-\frac{\pi|\alpha|}{96N^{2}}+{\cal O}\left(\frac{\pi|\alpha|^{3}}{N^{4}}\right)\ .$ (95)
@@
$\mathfrak{b}(\alpha)=\int_{0}^{\infty}\frac{dt}{\pi}\frac{\alpha}{t^{2}+\alpha^{2}}\ln\frac{1-e^{-t}}{1+e^{-t}}\equiv\ln\sqrt{\alpha}+\ln\frac{\Gamma\left(\frac{\alpha}{2\pi}+\frac{1}{2}\right)}{\sqrt{2\pi}\Gamma(\frac{\alpha}{2\pi}+1)}\ ,\ \alpha>0\ .$ (96)
@@ 2.2 The Wick contraction
${\cal M}_{2k+1,2l}(\beta_{i},z_{j})={\cal M}_{2k+1,2l}=(-1)^{l}\langle\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}\rangle\ .$ (165)
The extra factor of $(-1)^{l}$ is equivalent to the redefinition
@@
$(-1)^{l}\langle\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}\rangle={\cal M}_{2k+1,2l}(\beta_{i},z_{j})$
@@ 2.3 The vacuum overlap
where $A$ is the Glaisher constant. Eq. (206) will be shown in the Appendix. B. On the other hand, the term with $\tanh^{2}\frac{t}{4N}$ represents the sub-leading corrections and can be easily expanded to an arbitrary order in $\frac{1}{N}$
@@
$\sum_{i=1}^{n}\ln i!=(n+1)\ln n!-\ln\prod_{i=1}^{n}i^{i}=(n+1)\ln n!-\ln H(n)\ ,$ (210)
$\ln H(n)\rightarrow\left(\frac{n^{2}+n}{2}+\frac{1}{12}\right)\ln n-\frac{n^{2}}{4}+\ln A\ .$ (211)
@@ 2.4 Scaling limit of the two-point correlator
This is exactly the standard CFT two-point correlator on an infinite Cylinder. As such, after multiplying back the factor $|\langle\Omega_{R}|\Omega_{NS}\rangle|^{2}$, at small $r\ll L$, one has then
@@
It should be note that it is also possible to locate the two spin operators at different spatial positions, for example, one at $(\tau,i)$ and another at $(0,0)$, where $0\leq i\leq N-1$ is the spatial position in the lattice unit. Then the scaling limit in the spatial direction amounts to
$i\rightarrow\infty,\ N\rightarrow\infty,\ \frac{i}{N}\equiv\frac{x}{L}={\cal O}(1)\ .$ (230)
In this situation, it is not hard to show that one can introduce
$w=e^{-\frac{2\pi(r+ix)}{L}},\ \bar{w}=e^{-\frac{2\pi(r-ix)}{L}}\ ,$ (231)
and generalizes the Fredholm determinant formula to the following factorized form
$\langle\Omega_{NS}|\hat{\sigma}(r,x)\hat{\sigma}(0)|\Omega_{NS}\rangle=e^{-\frac{\pi r}{4L}}\det(1+{\cal K}_{ij}(w))\det(1+{\cal K}_{ij}(\bar{w}))\ .$ (232)
This is again consistent with the standard CFT formula DiFrancesco:1997nk.
@@ 3 Comparison with literature and further comments
$\kappa(\mu L,\mu L)=\frac{1}{\pi}\int_{1}^{\infty}\frac{dt}{t\sqrt{t^{2}-1}}\ln\frac{1-e^{-\mu Lt}}{1+e^{-\mu Lt}}=\frac{1}{2}\ln\mu L+{o}(1)\ ,$ (238)
@@
again consistent with the $d(1)=1$. Two particle functions clearly simplify to the $\Pi$ factors in Eq. (104) and Eq. (105). Moreover, in Appendix. A, we show that normalization function $\bar{\sigma}S(mL)$ of Fonseca:2001dc has the following small $m$ expansion
$\bar{\sigma}S(mL)=\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}\exp\bigg(\frac{\ln 2}{2\pi}mL-\frac{\ln^{2}2}{4\pi^{2}}(mL)^{2}-\frac{\zeta_{3}}{16\pi^{3}}(mL)^{3}+{\cal O}((mL)^{4})\bigg)\ .$ (240)
Given this, we conclude that the massless limit of the massive matrix elements in Fonseca:2001dc, up to overall phase factors, exactly reduce to $\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}$ multiplying our Eq. (104) and Eq. (105) for $\hat{\sigma}$. The normalization of Fonseca:2001dc ensures the small-$r$ limit of the massless two point correlator to be $r^{-\frac{1}{4}}$ without extra factors.
@@
where we have used the relation p(1)q(1)=2ip(1)q(1)=-2i in Eq. (71). As such, all the form factor squared in Gehlen_2008; Iorgov:2010yv, in the proper $g\rightarrow 1^{+}$ limit approached from the disordered side, agree with the results derived directly at the critical point.
@@
where $K(w)$ is the complete Elliptic integral of the second Kind. The $\frac{K(w)}{\pi}$ is exactly half of the one particle form factor contribution $f^{(1)}_{0,0}(t)$ to the diagonal correlator at zero separation Boukraa:2006bt. If such relations could be established at high orders, then the identity Eq. (249) becomes understandable: $\sigma_{00}^{2}=1$, while the form factor expansion in Boukraa:2006bt contains an overall $(1-t)^{\frac{1}{4}}$. This implies that all the form factors at zero separation must sum to $(1-t)^{-\frac{1}{4}}$. It is also possible that differential equations in $w$ could be established, similar to the Painleve VI for the parameter derivatives 10.3792/pjaa.56.405. Finally, it is also interesting to see how the form factors can be summed to the more general toroidal correlators at a finite temperature.
@@ Acknowledgements.
The author thanks Zoltán Bajnok for constructive discussions and introduction to the references Yurov:1991my; Fonseca:2001dc.
Appendix A The $s\rightarrow 0^{+}$ expansion of the crossover function
In this appendix, we compute analytically the small $s\rightarrow 0^{+}$ expansion of the function $g_{I}(s)$ of Sachdev:1995bf. We start from the definition with $s>0$
$-f_{I}(s)=\frac{1}{\pi}\int_{0}^{\infty}dy\ln\frac{1-e^{-\sqrt{y^{2}+s^{2}}}}{1+e^{-\sqrt{y^{2}+s^{2}}}}\ ,$ (251)
$\ln g_{I}(s)=\int_{s}^{1}\frac{ds^{\prime}}{s^{\prime}}\left((f^{\prime}_{I}(s^{\prime}))^{2}-\frac{1}{4}\right)+\int_{1}^{\infty}\frac{ds^{\prime}}{s^{\prime}}(f_{I}^{\prime}(s))^{2}\ .$ (252)
To perform the integral, it is convenient to use the following Mellin-Barnes representation of $f_{I}^{\prime}(s)$
$-f^{\prime}_{I}(s)=\int_{{\rm Re}(u)>0}\frac{du}{2\pi i}s^{-u}{\cal M}(u)\ ,$ (253)
${\cal M}(u)=\frac{\left(2^{u+1}-1\right)\Gamma^{2}\left(\frac{u+1}{2}\right)\zeta(u+1)}{2\pi}\ .$ (254)
The above can be established by expanding the logarithm into the Taylor series
$\ln\frac{1-e^{-\sqrt{y^{2}+s^{2}}}}{1+e^{-\sqrt{y^{2}+s^{2}}}}=-2\sum_{k=0}^{\infty}\frac{e^{-(2k+1)\sqrt{y^{2}+s^{2}}}}{2k+1}\ ,$ (255)
and Mellin-transform the individual terms. It is easy to see that
${\rm Res}\bigg({\cal M}(u)\bigg)(u=0)=\frac{1}{2}\ ,$ (256)
consistent with the subtraction term $\frac{1}{4}$. Given the above, it is then straightforward to establish the following double Barnes integral
$\ln g_{I}(s)=\int\int_{{{\rm Re}(u)-1}<{\rm Re}(z)<0<{\rm Re}(u)<1}\frac{dzdu}{(2\pi i)^{2}}\frac{s^{-z}-1}{z}{\cal M}(u){\cal M}(z-u)$
$+\frac{1}{2}\int_{-1<\rm Re(u)<0}\frac{du}{2\pi i}\frac{s^{-u}-1}{u}{\cal M}(u)+\int\int_{0<{\rm Re}(u)<{\rm Re}(z)<1}\frac{dzdu}{(2\pi i)^{2}}\frac{{\cal M}(u){\cal M}(z-u)}{z}\ .$ (257)
Here, the first two terms are due to the integral in the $(0,1)$ region, while the last term is due to the integral in the $(1,\infty)$ region. The $\frac{1}{4}$ subtraction is achieved by appropriate contour shifting. We now shift the $z$ contour in the last term to the ${\rm Re}(z)<0$ region to cancel the $s$ independent terms. There are $z=0$ and $z=u$ two poles. Using the fact that
${\rm Res}\left(\frac{{\cal M}(u)}{u}\right)(u=0)=0\ ,$ (258)
one finally obtains
$\ln g_{I}(s)=\int\int_{{{\rm Re}(u)-1}<{\rm Re}(z)<0<{\rm Re}(u)<1}\frac{dzdu}{(2\pi i)^{2}}\frac{s^{-z}}{z}{\cal M}(u){\cal M}(z-u)$
$+\frac{1}{2}\int_{-1<\rm Re(u)<0}\frac{du}{2\pi i}\frac{s^{-u}}{u}{\cal M}(u)+\int_{0<{\rm Re}(u)<1}\frac{du}{2\pi i}{\cal M}(u){\cal M}(-u)\ .$ (259)
Now, using the functional equation of the $\zeta$ function, it is straightforward to show that
${\cal M}(u){\cal M}(-u)\equiv\left(2-2^{u}\right)\left(2-2^{-u}\right)\Gamma(u)\Gamma(-u)\zeta(u)\zeta(-u)\ .$ (260)
As such, the last term of Eq. (259) exactly leads to the Barnes-$G$ product using the Barnes integral Eq. (212). On the other hand, in the first two terms, one can shift the contours to the left to perform systematically the small $s$ expansion. The results of the expansion up to the third order read
$\ln g_{I}(s)\bigg|_{s\rightarrow 0^{+}}=\ln\bigg(\pi^{\frac{1}{4}}G\left(\frac{1}{2}\right)G\left(\frac{3}{2}\right)\bigg)+\frac{\ln 2}{\pi}s-\frac{\ln^{2}2}{2\pi^{2}}s^{2}-\frac{\zeta_{3}}{8\pi^{3}}s^{3}+{\cal O}(s^{4})\ .$ (261)
In particular, the leading term exactly reproduces the constant ${\cal C}_{1}$ in Sachdev:1995bf extracted numerically. At higher orders, there will be $\zeta_{2k+1}\zeta_{2l+1}$ type terms. The expansion has a convergence radius equals to $\pi$.
Given the expansion of $\ln g_{I}(s)$, the expansion of the function $S(mL)$ in Fonseca:2001dc can also be found using the relation
$\ln S(mL)=\frac{1}{2}\ln g_{I}\left(mL\right)-\frac{1}{8}\ln mL\ ,$ (262)
which is equivalent to Eq. (A13) of Fonseca:2001dc. Especially, using the expansion Eq. (261), in the massless limit the $\bar{\sigma}m^{\frac{1}{8}}S(mL)$ is finite
$\lim_{m\rightarrow 0}m^{\frac{1}{8}}2^{\frac{1}{12}}e^{-\frac{1}{8}}A^{\frac{3}{2}}S(mL)=\frac{2^{\frac{1}{12}}e^{-\frac{1}{8}}A^{\frac{3}{2}}}{L^{\frac{1}{8}}}(\pi e)^{\frac{1}{8}}2^{\frac{1}{24}}A^{-\frac{3}{2}}=\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}\ .$ (263)
Using the Eq. (2.15) of Fonseca:2001dc and the massless limits of their leg functions discussed in Sec. 3, this implies that the overall normalization of the spin operator in Fonseca:2001dc is exactly $\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}$ times our $\hat{\sigma}$. In particular, at small $r$, the two point correlator approaches a simple power-law $r^{-\frac{1}{4}}$ without extra factors.
Appendix B The integral Eq. (206)
In this appendix, we prove the integral formula Eq. (206). We first change the variable from $\sqrt{t}\rightarrow t$, and introduce another integral representation for $(1-t)(\ln t)^{-1}$, leading to
$I_{0}=\frac{1}{2}\int_{0}^{1}\frac{dt}{\ln t}\frac{1-t}{(1+t)^{2}}=-\frac{1}{2}\int_{0}^{1}dx\int_{0}^{1}dt\frac{t^{x}}{(1+t)^{2}}\ .$ (264)
It convergences absolutely. The $t$ integral can be integrated to a special case of the Gauss Hypergeometric function and can be expressed in terms of the digamma function, leading to
$\int_{0}^{1}dt\frac{t^{x}}{(1+t)^{2}}=\frac{1}{2}+\frac{x}{2}\left(\psi\left(\frac{1+x}{2}\right)-\psi\left(1+\frac{x}{2}\right)\right)\ ,$ (265)
$I_{0}=-\frac{1}{4}-\frac{1}{4}\int_{0}^{1}dxx\left(\psi\left(\frac{1+x}{2}\right)-\psi\left(1+\frac{x}{2}\right)\right)\ .$ (266)
The integral here is similar to that of the Ising connecting computation Tracy:1990tn. Using $\psi(x)=(\ln\Gamma(x))^{\prime}$ and partial integrating as in Tracy:1990tn, one has
$I_{0}=\frac{1}{2}\ln\frac{\Gamma\left(\frac{1}{2}\right)}{\Gamma\left(1\right)}+\ln\frac{\Gamma_{2}^{2}(1)}{\Gamma_{2}\left(\frac{1}{2}\right)\Gamma_{2}\left(\frac{3}{2}\right)}=\ln\bigg(\pi^{\frac{1}{4}}G\left(\frac{1}{2}\right)G\left(\frac{3}{2}\right)\bigg)\ .$ (267)
Here we have used the relation $G({\cal Z})=\frac{1}{\Gamma_{2}({\cal Z})}$ between the Barnes $G$ function and the double gamma function $\Gamma_{2}({\cal Z})$. This is exactly Eq. (206).